OpenStudy (mertsj):

@mr. math Check that implicit differentiation problem of the circle x^2+y^2=16. The derivative of a constant is 0

OpenStudy (mr.math):

I just posted my comment there.

OpenStudy (anonymous):

\[2x + 2y\frac{dy}{dx} = 0\] \[\frac{dy}{dx} = -\frac{x}{y}\]